Find the transfer function of the system, R=20 C=0.5f Cy(t)
Answer #1

for Given Cúcuit bansfer function R=ae (= 0.5€ Writing the dicterential Equation with the U kirchofts haw help of - (+) : Rior Vis] = SRC465) + 4(s) * v(s) = 4(s) [it SR] - § ys) - V(s) IHSRC f in Putting V Value above of R= de O Grauoker and C= 0·5

By using the laplace transformation of interger to 1/s ,here we can easilly find the voltage across capacitor ,go through the image for step by step solution.